# Adversarial Policy Optimization for Offline Preference-based Reinforcement Learning

### v1

@@ 1.1 Related Work
Provable Offline PbRL. While there has been a growing amount of research on online PbRL, the theoretical understanding of offline PbRL remains relatively limited. A primary challenge in offline PbRL, much like in offline standard RL, is ensuring sufficient conservatism in the model. Zhu et al. (2023) addressed this challenge by proposing a pessimistic maximum likelihood estimation (MLE) algorithm for the linear model with known transitions. Zhan et al. (2024a) extended this idea to general function approximation, highlighting the importance of trajectory concentrability in establishing a lower bound for sample complexity. Despite the provable sample complexity bound, their proposed algorithm, FREEHAND-transition, relies on solving $\operatorname*{arg\,max}_{\pi}\operatorname*{arg\,min}_{r\in\hat{\mathcal{R}}}\operatorname*{arg\,min}_{P\in\hat{\mathcal{P}}}\{\mathbb{E}_{\tau\sim P,\pi}[r(\tau)]-\mathbb{E}_{\tau\sim P^{\star},\pi}[r(\tau)]\}$ where $\hat{\mathcal{R}}$ is the confidence set of rewards and $\hat{\mathcal{P}}$ is the confidence set of transitions, which is intractable in practice. Pace et al. (2024) introduced an algorithm achieving conservatism through explicit uncertainty penalties defined as $u_{R}(\tau)=\sup_{r_{1},r_{2}\in\hat{\mathcal{R}}}|r_{1}(\tau)-r_{2}(\tau)|$ (reward uncertainty) and $u_{P}(s,a)=\sup_{P_{1},P_{2}\in\hat{\mathcal{P}}}\left\|P_{1}(\cdot\mid s,a)-P_{2}(\cdot\mid s,a)\right\|_{1}$ (transition uncertainty). Even evaluating this function is intractable with general function approximation, but Pace et al. (2024) requires optimizing $\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\tau\sim\hat{P},\pi}[\hat{r}(\tau)-u_{R}(\tau)-u_{P}(\tau)]$. Chang et al. (2024) explored a slightly different scenario where the data collection policy is known and online interaction is allowed. They demonstrated that a simple natural policy gradient combined with MLE reward is provably efficient, but their sample complexity bound is affected by an additional concentrability coefficient relative to KL-regularized policies.
@@ 2 Preliminaries
where we denote $r^{\star}(\tau)=\sum^{H}_{h=1}r^{\star}_{h}(s_{h},a_{h})$ for given trajectory $\tau=(s_{1},a_{1},\dots,s_{H},a_{H})$. Additionally, we assume that $\kappa=1/(\inf_{x\in[-R,R]}\Phi^{\prime}(x))$, where $R$ is a bound on trajectory returns, is finite. When $\Phi$ is set to be the sigmoid function $\sigma(x)=1/(1+\exp(-x))$, we obtain the widely used Bradely-Terry-Luce model (Bradley & Terry, 1952). In addition to the preference dataset, we have an unlabeled trajectory dataset $\mathcal{D}_{\text{traj}}=\{(\tau^{0,n},\tau^{1,n})\}^{N}_{n=1}$ where the trajectory pairs are sampled i.i.d. by executing the reference policy $\mu$. The agent’s goal is to find an $\epsilon$-optimal policy $\hat{\pi}$ with respect to the optimal policy $\pi^{\star}$, which satisfies $V^{\pi^{\star}}_{1,r^{\star}}(s_{1})-V^{\hat{\pi}}_{1,r^{\star}}(s_{1})\leq\epsilon$.
@@
^$\hat{\mathcal{L}}_{R}(r)=-\underset{(\tau^{0},\tau^{1},y)\sim\mathcal{D}_{\text{pref}}}{\mathbb{E}}\left[\mathbbm{1}\left\{y=1\right\}\cdot\log\Phi(r(\tau^{1})-r(\tau^{0}))+\mathbbm{1}\left\{y=0\right\}\cdot\log\Phi(r(\tau^{0})-r(\tau^{1}))\right].$
@@ 3.2 Adversarial Optimization for PbRL
1:Input: Number of rollouts $K_{1},K_{2}$, KL regularization $\eta$, $\pi^{1}_{h}=\text{Unif}(\mathcal{A})$ for all $h\in[H]$
@@ 3.3 APPO: Reparameterized Algorithm for Unknown Transition
1:Input: KL regularization $\eta$, Initial policy $\pi^{1}_{h}=\text{Unif}(\mathcal{A})$ for all $h\in[H]$
@@
^$\hat{\mathcal{E}}_{\mathcal{D}_{\text{traj}}}(f;\hat{P},\hat{r}):=\mathbb{E}_{(\tau_{0},\tau_{1})\sim\mu}\left[\left|\{r^{\pi^{t}}_{\hat{P},f}(\tau^{0})-r^{\pi^{t}}_{\hat{P},f}(\tau^{1})\}-\{\hat{r}(\tau^{0})-\hat{r}(\tau^{1})\}\right|\right]\approx\mathcal{E}(f;P^{\star},\hat{r}),$
@@ 4 Theoretical Analysis
The bounded $C_{\text{TR}}$ ensures that the support of $d^{\mu}$ sufficiently covers the support of $d^{\pi^{\star}}$, similar to the concentrability condition in Zhan et al. (2024a). As a result, we expect $\mathcal{D}_{\text{traj}}$ to contain high-quality trajectories. The lower bound in Zhan et al. (2024a) shows that the trajectory concentrability is essential in offline PbRL. This implies that offline PbRL is strictly more challenging than offline standard RL, where step-wise concentrability is sufficient to achieve a performance guarantee (Uehara & Sun, 2022). We now present the sample complexity bound.
@@
Discussion on Theorem 4.1. Our analysis naturally extends to infinite function classes by applying the standard covering number argument, replacing the cardinalities $|\mathcal{R}|$, $|\mathcal{P}|$, and $|\mathcal{F}|$ with covering numbers. To our knowledge, FREEHAND-transition Zhan et al. (2024a) and Sim-OPRL Pace et al. (2024) are the only statistically efficient algorithms for offline PbRL in stochastic MDPs. Our sample complexity bound matches theirs for labeled data ($M$). However, FREEHAND-transition and Sim-OPRL require $\Theta\left(\frac{C_{P}^{2}R^{2}H^{2}\log(H|\mathcal{P}|/\delta)}{\epsilon^{2}}\right)$ unlabeled trajectories where $C_{P}$ is the trajectory concentrability for transition 222 Zhan et al. (2024a) consider reward functions defined over trajectories, thus their reward class $\mathcal{G}_{r}$ is comparable to our $\mathcal{R}^{H}$. Pace et al. (2024) assume a homogeneous reward, so their presented bound is tighter by $H$. They use bracketing numbers in their bound, but we write here $|\mathcal{P}|$ for simplicity.. This highlights a trade-off: While FREEHAND-transition and Sim-OPRL have tighter bounds for unlabeled data ($N$), they are computationally intractable. FREEHAND-transition requires solving a nearly intractable nested optimization problem, while Sim-OPRL relies on an uncertainty penalty defined by the width of confidence sets. Therefore, our APPO is the first offline PbRL algorithm to achieve both provable statistical efficiency and computational efficiency.
@@ 5 Practical Implementation of APPO
BPT-500 $88.33$±$\pm 4.76$ $10.08$±$\pm 7.57$ $22.87$±$\pm 9.06$ $3.93$±$\pm 4.34$ $34.73$±$\pm 13.9$ $\bm{53.52}$±$\pm 13.9$
box-close-500 $93.40$±$\pm 3.10$ $\bm{29.12}$±$\pm 13.2$ $0.33$±$\pm 1.16$ $10.20$±$\pm 11.5$ $5.93$±$\pm 5.81$ $\bm{18.24}$±$\pm 15.6$
dial-turn-500 $\pm 5.47$ $61.44$±$\pm 6.08$ $\bm{68.67}$±$\pm 12.4$ $26.67$±$\pm 22.2$ $31.53$±$\pm 12.5$ $\bm{80.96}$±$\pm 4.49$
sweep-500 $98.33$±$\pm 1.87$ $\bm{86.96}$±$\pm 6.93$ $43.07$±$\pm 24.6$ $10.47$±$\pm 15.8$ $27.20$±$\pm 23.8$ $26.80$±$\pm 5.32$
BPT-wall-500 $56.27$±$\pm 6.32$ $0.32$±$\pm 0.30$ $0.87$±$\pm 1.43$ $0.80$±$\pm 1.51$ $8.93$±$\pm 9.84$ $\bm{64.32}$±$\pm 21.0$
sweep-into-500 $78.80$±$\pm 7.96$ $\bm{28.40}$±$\pm 5.47$ $20.53$±$\pm 8.26$ $23.07$±$\pm 7.02$ $\bm{32.20}$±$\pm 7.35$ $24.08$±$\pm 5.91$
drawer-open-500 $100.00$±$\pm 0.00$ $\bm{98.00}$±$\pm 2.32$ $88.73$±$\pm 11.6$ $35.93$±$\pm 11.2$ $19.00$±$\pm 13.6$ $87.68$±$\pm 10.0$
lever-pull-500 $98.47$±$\pm 1.77$ $\bm{79.28}$±$\pm 2.95$ $\bm{82.40}$±$\pm 22.7$ $10.13$±$\pm 12.2$ $31.20$±$\pm 15.8$ $\bm{75.76}$±$\pm 7.17$
BPT-1000 $88.33$±$\pm 4.76$ $8.48$±$\pm 5.80$ $18.27$±$\pm 10.6$ $3.20$±$\pm 3.04$ $36.67$±$\pm 17.4$ $\bm{59.04}$±$\pm 19.0$
box-close-1000 $93.40$±$\pm 3.10$ $\bm{27.04}$±$\pm 14.5$ $2.27$±$\pm 2.86$ $9.33$±$\pm 9.60$ $6.73$±$\pm 8.41$ $\bm{34.24}$±$\pm 18.5$
dial-turn-1000 $\pm 5.47$ $69.44$±$\pm 4.70$ $68.80$±$\pm 5.50$ $36.40$±$\pm 21.9$ $43.93$±$\pm 13.4$ $\bm{81.44}$±$\pm 6.73$
sweep-1000 $98.33$±$\pm 1.87$ $\bm{87.52}$±$\pm 7.87$ $29.13$±$\pm 14.6$ $8.73$±$\pm 16.4$ $38.33$±$\pm 24.9$ $17.36$±$\pm 12.4$
BPT-wall-1000 $56.27$±$\pm 6.32$ $0.48$±$\pm 0.47$ $2.13$±$\pm 2.96$ $0.27$±$\pm 0.85$ $14.07$±$\pm 11.5$ $\bm{62.96}$±$\pm 18.4$
sweep-into-1000 $78.80$±$\pm 7.96$ $\bm{26.00}$±$\pm 5.53$ $20.27$±$\pm 7.84$ $\bm{23.33}$±$\pm 7.80$ $\bm{30.40}$±$\pm 7.74$ $18.16$±$\pm 11.1$
drawer-open-1000 $100.00$±$\pm 0.00$ $\bm{98.40}$±$\pm 2.82$ $\bm{95.40}$±$\pm 7.27$ $36.47$±$\pm 7.30$ $28.53$±$\pm 18.4$ $\bm{98.56}$±$\pm 2.68$
lever-pull-1000 $98.47$±$\pm 1.77$ $\bm{88.96}$±$\pm 3.94$ $72.93$±$\pm 10.2$ $8.53$±$\pm 9.96$ $40.40$±$\pm 17.4$ $76.96$±$\pm 4.40$
Average Rank - $2.316$ $3.125$ $4.375$ $3.063$ $\bm{2.125}$

### v2

@@ 1.1 Related Work
Provable Offline PbRL. While there has been a growing amount of research on online PbRL, the theoretical understanding of offline PbRL remains relatively limited. A primary challenge in offline PbRL, much like in offline standard RL, is ensuring sufficient conservatism in the model. Zhu et al. (2023) addressed this challenge by proposing a pessimistic maximum likelihood estimation (MLE) algorithm for the linear model with known transitions. Zhan et al. (2024a) extended this idea to general function approximation, highlighting the importance of trajectory concentrability in establishing a lower bound for sample complexity. Despite the provable sample complexity bound, their proposed algorithm, FREEHAND-transition, relies on solving $\operatorname*{arg\,max}_{\pi}\operatorname*{arg\,min}_{r\in\hat{\mathcal{R}}}\operatorname*{arg\,min}_{P\in\hat{\mathcal{P}}}\{\mathbb{E}_{\tau\sim P,\pi}[r(\tau)]-\mathbb{E}_{\tau\sim P^{\star},\pi}[r(\tau)]\}$ where $\hat{\mathcal{R}}$ is the confidence set of rewards and $\hat{\mathcal{P}}$ is the confidence set of transitions, which is intractable in practice. Pace et al. (2024) study a different but related setting, where the agent elicits high-quality preference information from offline data. Their method achieves conservatism through explicit uncertainty penalties defined as $u_{R}(\tau)=\sup_{r_{1},r_{2}\in\hat{\mathcal{R}}}|r_{1}(\tau)-r_{2}(\tau)|$ (reward uncertainty) and $u_{P}(s,a)=\sup_{P_{1},P_{2}\in\hat{\mathcal{P}}}\left\|P_{1}(\cdot\mid s,a)-P_{2}(\cdot\mid s,a)\right\|_{1}$ (transition uncertainty), which makes it intractable with general function approximation. Chang et al. (2024) also explored a slightly different scenario where the data collection policy is known and online interaction is allowed. They demonstrated that a simple natural policy gradient combined with MLE reward is provably efficient, but their sample complexity bound is affected by an additional concentrability coefficient relative to KL-regularized policies.
@@ 2 Preliminaries
where we denote $r^{\star}(\tau)=\sum^{H}_{h=1}r^{\star}_{h}(s_{h},a_{h})$ for given trajectory $\tau=(s_{1},a_{1},\dots,s_{H},a_{H})$. Additionally, we assume that $\kappa=1/(\inf_{x\in[-R,R]}\Phi^{\prime}(x))$, where $R$ is a bound on trajectory returns, is finite. When $\Phi$ is set to be the sigmoid function $\sigma(x)=1/(1+\exp(-x))$, we obtain the widely used Bradely-Terry-Luce model (Bradley & Terry, 1952). In addition to the preference dataset, we have an unlabeled trajectory dataset $\mathcal{D}_{\text{traj}}=\{(\tau^{n,0},\tau^{n,1})\}^{N}_{n=1}$ where the trajectory pairs are sampled i.i.d. by executing the reference policy $\mu$. The agent’s goal is to find an $\epsilon$-optimal policy $\hat{\pi}$ with respect to the optimal policy $\pi^{\star}$, which satisfies $V^{\pi^{\star}}_{1,r^{\star}}(s_{1})-V^{\hat{\pi}}_{1,r^{\star}}(s_{1})\leq\epsilon$.
@@
^$\hat{\mathcal{L}}_{R}(r)=-\underset{(\tau^{0},\tau^{1},y)\sim\mathcal{D}_{\text{pref}}}{\mathbb{E}}\left[y\cdot\log\Phi(r(\tau^{1})-r(\tau^{0}))+(1-y)\cdot\log\Phi(r(\tau^{0})-r(\tau^{1}))\right].$
@@ 3.2 Adversarial Optimization for PbRL
1:Input: Number of rollouts $K_{1},K_{2}$, constant $\eta$, $\pi^{1}_{h}=\text{Unif}(\mathcal{A})$ for all $h\in[H]$
@@ 3.3 APPO: Reparameterized Algorithm for Unknown Transition
1:Input: Constant $\eta$, Initial policy $\pi^{1}_{h}=\text{Unif}(\mathcal{A})$ for all $h\in[H]$
@@
^$\hat{\mathcal{E}}_{\mathcal{D}_{\text{traj}}}(f;\hat{P},\hat{r}):=\mathbb{E}_{(\tau_{0},\tau_{1})\sim\mathcal{D}_{\text{traj}}}\left[\left|\{r^{\pi^{t}}_{\hat{P},f}(\tau^{0})-r^{\pi^{t}}_{\hat{P},f}(\tau^{1})\}-\{\hat{r}(\tau^{0})-\hat{r}(\tau^{1})\}\right|\right]\approx\mathcal{E}(f;P^{\star},\hat{r}),$
@@ 4 Theoretical Analysis
The bounded $C_{\text{TR}}$ ensures that the support of $d^{\mu}$ sufficiently covers the support of $d^{\pi^{\star}}$, similar to the concentrability condition in Zhan et al. (2024a) [^2] Our analysis remains valid under an alternative definition based on the reward model error ratio, similar to that used by Zhan et al. (2024a).. As a result, we expect $\mathcal{D}_{\text{traj}}$ to contain high-quality trajectories. The lower bound in Zhan et al. (2024a) shows that the trajectory concentrability is essential in offline PbRL. This implies that offline PbRL is strictly more challenging than offline standard RL, where step-wise concentrability is sufficient to achieve a performance guarantee (Uehara & Sun, 2022). We now present the sample complexity bound.
@@
Discussion on Theorem 4.1. Our analysis naturally extends to infinite function classes by applying the standard covering number argument, replacing the cardinalities $|\mathcal{R}|$, $|\mathcal{P}|$, and $|\mathcal{F}|$ with covering numbers. To our knowledge, FREEHAND-transition Zhan et al. (2024a) is the only statistically efficient algorithm for offline PbRL in stochastic MDPs. Our sample complexity bound matches theirs for labeled data ($M$). However, FREEHAND-transition requires $\Theta\left(\frac{C_{P}^{2}R^{2}H^{2}\log(H|\mathcal{P}|/\delta)}{\epsilon^{2}}\right)$ unlabeled trajectories where $C_{P}$ is the trajectory concentrability for transition 333 Zhan et al. (2024a) consider reward functions defined over trajectories, thus their reward class $\mathcal{G}_{r}$ is comparable to our $\mathcal{R}^{H}$. They use bracketing numbers in their bound, but we write here $|\mathcal{P}|$ for simplicity.. This highlights a trade-off: While FREEHAND-transition has tighter bounds for unlabeled data ($N$), it is computationally intractable. That is, FREEHAND-transition requires solving a nearly intractable nested optimization problem. Therefore, our APPO is the first offline PbRL algorithm to achieve both provable statistical efficiency and computational efficiency.
@@ 5 Practical Implementation of APPO
BPT-500 88.33 ±$\pm 4.76$ 10.08 ±$\pm 7.57$ 22.87 ±$\pm 9.06$ 3.93 ±$\pm 4.34$ 34.73 ±$\pm 13.9$ 53.52±$\pm 13.9$
box-close-500 93.40 ±$\pm 3.10$ 29.12±$\pm 13.2$ 0.33 ±$\pm 1.16$ 10.20 ±$\pm 11.5$ 5.93 ±$\pm 5.81$ 18.24±$\pm 15.6$
dial-turn-500 75.40 ±$\pm 5.47$ 61.44 ±$\pm 6.08$ 68.67±$\pm 12.4$ 26.67 ±$\pm 22.2$ 31.53 ±$\pm 12.5$ 80.96±$\pm 4.49$
sweep-500 98.33 ±$\pm 1.87$ 86.96±$\pm 6.93$ 43.07 ±$\pm 24.6$ 10.47 ±$\pm 15.8$ 27.20 ±$\pm 23.8$ 26.80 ±$\pm 5.32$
BPT-wall-500 56.27 ±$\pm 6.32$ 0.32 ±$\pm 0.30$ 0.87 ±$\pm 1.43$ 0.80 ±$\pm 1.51$ 8.93 ±$\pm 9.84$ 64.32±$\pm 21.0$
sweep-into-500 78.80 ±$\pm 7.96$ 28.40±$\pm 5.47$ 20.53 ±$\pm 8.26$ 23.07 ±$\pm 7.02$ 32.20±$\pm 7.35$ 24.08 ±$\pm 5.91$
drawer-open-500 100.00 ±$\pm 0.00$ 98.00±$\pm 2.32$ 88.73 ±$\pm 11.6$ 35.93 ±$\pm 11.2$ 19.00 ±$\pm 13.6$ 87.68 ±$\pm 10.0$
lever-pull-500 98.47 ±$\pm 1.77$ 79.28±$\pm 2.95$ 82.40±$\pm 22.7$ 10.13 ±$\pm 12.2$ 31.20 ±$\pm 15.8$ 75.76±$\pm 7.17$
BPT-1000 88.33 ±$\pm 4.76$ 8.48 ±$\pm 5.80$ 18.27 ±$\pm 10.6$ 3.20 ±$\pm 3.04$ 36.67 ±$\pm 17.4$ 59.04±$\pm 19.0$
box-close-1000 93.40 ±$\pm 3.10$ 27.04±$\pm 14.5$ 2.27 ±$\pm 2.86$ 9.33 ±$\pm 9.60$ 6.73 ±$\pm 8.41$ 34.24±$\pm 18.5$
dial-turn-1000 75.40 ±$\pm 5.47$ 69.44 ±$\pm 4.70$ 68.80 ±$\pm 5.50$ 36.40 ±$\pm 21.9$ 43.93 ±$\pm 13.4$ 81.44±$\pm 6.73$
sweep-1000 98.33 ±$\pm 1.87$ 87.52±$\pm 7.87$ 29.13 ±$\pm 14.6$ 8.73 ±$\pm 16.4$ 38.33 ±$\pm 24.9$ 17.36 ±$\pm 12.4$
BPT-wall-1000 56.27 ±$\pm 6.32$ 0.48 ±$\pm 0.47$ 2.13 ±$\pm 2.96$ 0.27 ±$\pm 0.85$ 14.07 ±$\pm 11.5$ 62.96±$\pm 18.4$
sweep-into-1000 78.80 ±$\pm 7.96$ 26.00±$\pm 5.53$ 20.27 ±$\pm 7.84$ 23.33±$\pm 7.80$ 30.40±$\pm 7.74$ 18.16 ±$\pm 11.1$
drawer-open-1000 100.00 ±$\pm 0.00$ 98.40±$\pm 2.82$ 95.40±$\pm 7.27$ 36.47 ±$\pm 7.30$ 28.53 ±$\pm 18.4$ 98.56±$\pm 2.68$
lever-pull-1000 98.47 ±$\pm 1.77$ 88.96±$\pm 3.94$ 72.93 ±$\pm 10.2$ 8.53 ±$\pm 9.96$ 40.40 ±$\pm 17.4$ 76.96 ±$\pm 4.40$
Average Rank - 2.316 3.125 4.375 3.063 2.125